# Existence of attracting periodic orbits in 3-dimensional strongly 2-cooperative systems

### v1

@@ Abstract
The flow of a $k$-cooperative system maps the set of vectors with up to $(k-1)$ sign variations to itself. In particular, $1$-cooperative systems are just cooperative systems. Strongly $2$-cooperative systems satisfy a strong Poincaré-Bendixson property: any bounded solution that evolves in a compact set containing no equilibria converges to a periodic orbit. For $3$-dimensional strongly $2$-cooperative nonlinear systems, we provide a sufficient condition that guarantees the existence of an invariant compact set in the state space that includes an attracting periodic orbit. We show that our theoretical results unify and generalize known results on the existence of a periodic solution in two well-known models in biochemistry, a 3D Goodwin oscillator model and the 3D Field-Noyes ordinary-differential-equation (ODE) model for the Belousov-Zhabotinskii reaction, while simplifying the proofs.
Index Terms:
@@
I Introduction
@@
In cooperative systems, an increase in one of the state variables can never decrease the derivative of another state variable: the state variables “cooperate” with one another. The flow of such systems maps the non-negative orthant to itself. More precisely, the flow maps the set of vectors with zero sign variations to itself. Cooperative systems have many special asymptotic properties: Hirsch’s quasi-convergence theorem asserts that in a strongly cooperative system almost every bounded solution converges to an equilibrium point; also, cooperative systems cannot have an attracting periodic orbit [30]. Cooperative systems (as well as interconnections of cooperative systems) and their generalization to cooperative control systems [2] have found numerous applications in fields including systems biology and chemistry [3, 7, 8, 18, 21, 31], dynamic neural networks [29], and social dynamics [28].
@@
A recent generalization of nonlinear cooperative systems is called $k$-cooperative systems [35]. A system is $k$-cooperative if its flow maps the set of vectors with up to $k-1$ sign variations to itself. Just like cooperativity, $k$-cooperativity depends on the sign structure of the system Jacobian, and can be inferred even when the exact values of certain parameters are unknown. $1$-cooperative systems are cooperative systems, whereas $(n-1)$-cooperative systems (where $n$ is the dimension of the system) are, up to a coordinate transformation, competitive systems [35]. Another important case is strongly $2$-cooperative systems, as these satisfy a strong Poincaré-Bendixson property. This implies that if the closure of a trajectory of a bounded solution contains no equilibrium points, then that solution converges to a periodic orbit [35].
@@
It is interesting to identify cases where a bounded solution does not converge to an equilibrium, but to a periodic orbit, implying that the system admits at least one attracting periodic solution, thereby allowing for oscillatory behavior. For example, in the context of systems biology Goodwin [14] states that: “It is of fundamental importance to an understanding of cellular organization whether or not the dynamic activity of molecular control processes involves oscillatory behavior.” The analysis cannot be based on the theory of cooperative systems, because the existence of an attracting periodic orbit automatically implies that the system is not cooperative.
@@
Here, we consider 3-dimensional 2-cooperative systems. Our main result is a new simple sufficient condition for the existence of an attracting periodic orbit. The analysis is based on combining: (1) the strong Poincaré-Bendixson property of strongly $2$-cooperative systems; (2) the spectral properties of Jacobians of $2$-cooperative systems; and (3) an idea of Rauch [26] for constructing an invariant set $\Omega$ for a specific 3D model of a non-linear electric circuit such that $\Omega$ does not include an equilibrium.
We show that our theoretical results unify and generalize results on two specific non-linear 3D systems: a 3D Goodwin model and Field–Noyes 3D ODE model for the famous Belousov–Zhabotinskii reaction. The existence of a periodic solution is proven for the former model by Tyson [33] and for the latter model by Hastings and Murray [16]. However, in both cases, no information is provided on the stability of the periodic solution. Moreover, these papers rely on an application of the Schauder fixed-point theorem for proving existence of a periodic orbit, which is preceded by a fairly complicated preliminary analysis.
Our main result employs much simpler techniques than both [33] and [16] for proving existence of a periodic orbit and further implies that each of these systems admits at least one attracting periodic orbit. This is a fundamental extension of [33, 16] because, from an application viewpoint, a periodic orbit $\gamma$ which is not attracting is less interesting, as it yields no guarantees on oscillatory behavior, meaning that only solutions emanating from an initial condition precisely on the periodic orbit, i.e. $x(0)\in\gamma$, are guaranteed to induce oscillations.
We mention in passing that reassessing these papers from the point of view of $2$-cooperative systems shows that the proof in [16] is, in fact, incomplete. [^1] In particular, analyzing the sign structure of the sets $B_{i}$ in Eq. (20) in [16] and using the sign variation properties of 2-cooperative systems shows that the property described in this equation is wrong.
The remainder of this note is organized as follows. The next section reviews definitions and results that will be used subsequently. Section III details our main result. Section IV shows that our main result implies the existence of an attracting periodic orbit in a 3D Goodwin model and in the Field–Noyes 3D ODE model for the Belousov–Zhabotinskii reaction. The final section concludes, and describes several possible directions for further research.
II Notation and Preliminary Results
We denote vectors and matrices by lowercase and uppercase letters, respectively. For two vectors $x,y\in{\mathbb{R}}^{n}$, we write $x\geq y$ if $x_{i}\geq y_{i}$ for all $i$. The non-negative orthant in ${\mathbb{R}}^{n}$ is
${\mathbb{R}}^{n}_{\geq 0}:=\{x\in{\mathbb{R}}^{n}\,|\,x\geq 0\}.$
@@ II Notation and Preliminary Results
The transpose and determinant of a matrix $A$ are denoted by $A^{\top}$ and $\det(A)$ respectively, while $I_{n}$ denotes the $n\times n$ identity matrix. A square matrix $A$ is called Hurwitz if all its eigenvalues have a negative real part, unstable if it admits an eigenvalue with a positive real part, and Metzler if all its off-diagonal entries are non-negative. A matrix $\bar{A}$ is a sign matrix if every entry of $\bar{A}$ is either * (standing for “don’t care”), $\leq 0$, or $\geq 0$. A time-varying matrix $A(t)$ has the sign pattern $\bar{A}$ if the following three properties hold at all times $t$:
@@
    $a_{ij}(t)=0$ for all indices $i,j$ such that $\bar{a}_{ij}=0$;
@@
    $a_{ij}(t)\geq 0$ for all indices $i,j$ such that $\bar{a}_{ij}$ equals $\geq 0$;
@@
    $a_{ij}(t)\leq 0$ for all indices $i,j$ such that $\bar{a}_{ij}$ equals $\leq 0$,
while there is no restriction on $a_{ij}(t)$ when $\bar{a}_{ij}=*$. Given a set $S$, $\operatorname{int}(S)$ denotes its interior, and by $|S|$ its cardinality.
As mentioned earlier, the flow of a cooperative system maps ${\mathbb{R}}^{n}_{\geq 0}$ to ${\mathbb{R}}^{n}_{\geq 0}$, and it also maps ${\mathbb{R}}^{n}_{\leq 0}:=-({\mathbb{R}}^{n}_{\geq 0})$ to ${\mathbb{R}}^{n}_{\leq 0}$. In other words, the flow maps the set of vectors with zero sign variations to itself. A $k$-cooperative system maps the set of vectors with up to $k-1$ sign variations to itself.
Sign variations in a vector. Let $\sigma(x)$ denote the number of sign variations in a vector $x\in{\mathbb{R}}^{n}$ with no zero entries; for example, $\sigma\left(\begin{bmatrix}6.3&-\pi&1\end{bmatrix}^{\top}\right)=2$. The theory of totally positive matrices (see, e.g,. [9, 12, 25]) offers two useful generalizations of $\sigma(\cdot)$ to vectors that may include zero entries.
@@
Given $x\in{\mathbb{R}}^{n}\setminus\{0\}$, let $y$ denote the vector $x$ after deleting all its zero entries. Define
$s^{-}(x):=\sigma(y),$
and
$s^{+}(x):=\max_{z\in\mathcal{S}_{x}}\sigma(z),$
where $\mathcal{S}_{x}$ is the set of vectors obtained from $x$ by replacing each zero entry by either $+1$ or $-1$.
For example, for $x=\begin{bmatrix}-1&0&0&1&2\end{bmatrix}^{\top}$, we have $s^{-}(x)=\sigma(\begin{bmatrix}-1&1&2\end{bmatrix}^{\top})=1$, whereas $s^{+}(x)=\sigma(\begin{bmatrix}-1&1&-1&1&2\end{bmatrix}^{\top})=3$. If $x\in{\mathbb{R}}^{n}$ has no zero entries, then $s^{-}(x)=s^{+}(x)=\sigma(x)$. For the zero vector $0\in{\mathbb{R}}^{n}$, we define $s^{-}(0):=0$, and $s^{+}(0):=n-1$. Then
@@
Fix $x,y\in{\mathbb{R}}^{n}$, with $x\not=y$, so that $z:=x-y\not=0$; if $s^{+}(z)\leq 1$, then also the projection to the $2$-dimensional vector $\begin{bmatrix}z_{1}&z_{n}\end{bmatrix}^{\top}=\begin{bmatrix}x_{1}-y_{1}&x_{n}-y_{n}\end{bmatrix}^{\top}\not=0$.
@@
For example, $P^{1}_{-}={\mathbb{R}}^{n}_{\geq 0}\cup{\mathbb{R}}^{n}_{\leq 0}$, and $P^{1}_{+}=\operatorname{int}({\mathbb{R}}^{n}_{\geq 0}\cup{\mathbb{R}}^{n}_{\leq 0})$. More generally, it can be shown [35] that $P^{k}_{-}$ is closed, and that $P^{k}_{+}=\operatorname{int}(P^{k}_{-})$ for all $k$. By definition, these sets have a nested structure
$P^{1}_{-}\subset P^{2}_{-}\subset\dots\subset P^{n}_{-}={\mathbb{R}}^{n},$
$P^{1}_{+}\subset P^{2}_{+}\subset\dots\subset P^{n}_{+}={\mathbb{R}}^{n}.$
A set $C\subseteq{\mathbb{R}}^{n}$ is a cone if $x\in C$ implies that $\alpha x\in C$ for all $\alpha\in{\mathbb{R}}_{\geq 0}$. It is straightforward to verify that $P^{k}_{-}$ is a cone for any $k$. However, it is not a convex cone. For example, the vectors $y^{1}:=\begin{bmatrix}-2&1&1\end{bmatrix}^{\top}\in P^{2}_{-}$ and $y^{2}:=\begin{bmatrix}1&1&-2\end{bmatrix}^{\top}\in P^{2}_{-}$, but $\frac{1}{2}y^{1}+\frac{1}{2}y^{2}=\frac{1}{2}\begin{bmatrix}-1&2&-1\end{bmatrix}^{\top}\not\in P^{2}_{-}$.
To understand the geometry of $P^{k}_{-}$, recall that a closed set $C\subseteq{\mathbb{R}}^{n}$ is called a cone of rank $k$ if (1) $x\in C$ implies that $\alpha x\in C$ for all $\alpha\in{\mathbb{R}}$; and (2) $C$ contains a linear space of dimension $k$, and no linear space of a higher dimension [17]. For example, $P^{1}_{-}={\mathbb{R}}^{n}_{+}\cup{\mathbb{R}}^{n}_{-}$ is a cone of rank $1$. A cone $C$ of rank $k$ is called $k$-solid if there exists a $k$-dimensional linear subspace $V$ such that $V\setminus\{0\}\subseteq\operatorname{int}(C)$. For all $k\in\{1,\dots,n\}$, the set $P^{k}_{-}$ is a cone of rank $k$ that is $k$-solid [35].
Roughly speaking, if a dynamical system admits a $k$-solid cone as an invariant set, and if its trajectories can be projected in a one-to-one way on the $k$-dimensional linear subspace $V$ then its trajectories “behave” as those of a $k$-dimensional system. In particular, if $k=2$, then the trajectories “behave” as those of a planar dynamical system [27], thereby allowing to apply analysis techniques that are suitable for planar systems.
@@
Under the assumption that $t\to A(t)$ is continuous, the linear system (1) is $1$-positive (i.e., positive) iff $A(t)$ is Metzler for all $t$, that is, iff, for all $t$, $A(t)$ has the sign pattern
$\bar{A}_{1}:=\small\begin{bmatrix}*&\geq 0&\geq 0&\dots&\geq 0&\geq 0\\ \geq 0&*&\geq 0&\dots&\geq 0&\geq 0\\ \geq 0&\geq 0&*&\dots&\geq 0&\geq 0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ \geq 0&\geq 0&\geq 0&\dots&*&\geq 0\\ \geq 0&\geq 0&\geq 0&\dots&\geq 0&*\end{bmatrix}.$
The linear system (1) is $2$-positive iff, for all $t$$A(t)$ has the sign pattern
$\bar{A}_{2}:=\small\begin{bmatrix}*&\geq 0&0&\dots&0&\leq 0\\ \geq 0&*&\geq 0&\dots&0&0\\ 0&\geq 0&*&\dots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\dots&*&\geq 0\\ \leq 0&0&0&\dots&\geq 0&*\end{bmatrix},$ (2)
while it is strongly $2$-positive if, in addition, the matrix $A(t)$ is irreducible for almost all $t$ [35]. Since the upper-right and lower-left corner entries of $A(t)$ may take negative values, $A(t)$ is not necessarily Metzler.
Remark 1.
The system $\dot{x}(t)=\bar{A}_{2}x(t)$ may be interpreted as a multiagent system with agents $x_{1},\dots,x_{n}$ interconnected by bidirectional links with a ring topology. Each $x_{i}$, $i\in\{2,\dots,n-1\}$ receives arbitrary signed feedback from itself, and non-negative feedback from its left- and right-neighbours $x_{i-1}$ and $x_{i+1}$. Agent $x_{1}$ receives arbitrary signed feedback from itself, non-negative feedback from its right neighbour $x_{2}$ and non-positive feedback from its “cyclic left” neighbour $x_{n}$. Agent $x_{n}$ receives arbitrary signed feedback from itself, non-negative feedback from its left neighbour $x_{n-1}$ and non-positive feedback from its “cyclic right” neighbour $x_{1}$.
$k$-cooperative nonlinear dynamical systems. A nonlinear dynamical system is called $k$-cooperative if its associated variational system, which is a linear time-varying system, is $k$-positive for all times $t\geq 0$ and all points $x$ in the state space. Consider the time-invariant non-linear system
$\dot{x}=f(x),$ (3)
@@
$\dot{z}$ $=M(t)z,$ (4)
where
$M(t)$ $:=\int_{0}^{1}J\big{(}rx(t,a)+(1-r)x(t,b)\big{)}\mathop{}\!\mathrm{d}r.$
Note that (4) is an LTV, and that, if $J(x)$ satisfies some sign pattern condition for all $x\in\Omega$, then $M(t)$ has the same sign pattern for all $t\geq 0$, since sign pattern conditions are preserved under integration.
@@
For example, the system is $1$-cooperative if the variational system (4) is $1$-positive, that is, if $J(x)$ is Metzler for all $x$. Thus $1$-cooperativity is just cooperativity.
Remark 2.
Suppose that 00 is an equilibrium of (3). Fix $a\in\Omega$, with $a\not=0$, and define $z(t):=x(t,a)-x(t,0)=x(t,a)$. If the non-linear system (3) is strongly $k$-cooperative, it follows from Definitions 3 and 4 and from Eq. (4) that, if $s^{-}(x(t,a))\leq k-1$ for some time $t\geq 0$, then $s^{+}(x(\tau,a))\leq k-1$ for all $\tau>t$.
For an initial condition $a\in\Omega$, let $\omega(a)$ denote the omega limit set of the solution of (3) emanating from $x(0)=a$, namely, the set of all points $y\in\Omega$ for which there exists a sequence $(t_{n})_{n\in\mathbb{N}}$ in ${\mathbb{R}}$, with $\lim_{n\to\infty}t_{n}=+\infty$, such that $\lim_{n\to\infty}x(t_{n},a)=y$; see e.g. [32, p. 193].

### v2

# Convergence to periodic orbits in 3-dimensional strongly 2-cooperative systems

@@ Abstract
The flow of a $k$-cooperative system maps the set of vectors with up to $(k-1)$ sign variations to itself. Strongly $2$-cooperative systems satisfy a strong Poincaré-Bendixson property: any bounded solution that evolves in a compact set containing no equilibria converges to a periodic orbit. For $3$-dimensional strongly $2$-cooperative nonlinear systems, we provide a simple sufficient condition that guarantees the existence, in the state space, of an invariant compact set that includes no equilibrium points. Thus, any solution emanating from this set converges to a periodic orbit. We characterize explicitly the set of initial conditions from which the trajectory converges to a periodic solution. We demonstrate our theoretical results on two well-known models in biochemistry: a 3D Goodwin oscillator model and the 3D Field-Noyes ordinary-differential-equation (ODE) model for the Belousov-Zhabotinskii reaction.
keywords:
@@
thanks: R.K. and G.G. acknowledge support by the European Union through the ERC INSPIRE grant (project n. 101076926). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the EU, the European Research Executive Agency or the European Research Council. Neither the EU nor the granting authority can be held responsible for them.
1 Introduction
@@
In cooperative systems, an increase in one of the state variables can never decrease the derivative of another state variable: the state variables “cooperate” with one another. The flow of such systems maps the non-negative orthant to itself. Precisely, the flow maps the set of vectors with zero sign variations to itself. Cooperative systems have special asymptotic properties: they cannot have an attracting periodic orbit (Smith, 1995) and, in a strongly cooperative system, almost every bounded solution converges to an equilibrium point (Hirsch’s quasi-convergence theorem). Cooperative systems and their generalization to cooperative control systems (Angeli and Sontag, 2003) have found numerous applications in systems biology and chemistry (Angeli et al., 2004; Blanchini and Giordano, 2021; Donnell et al., 2009; Leenheer et al., 2007; Margaliot and Tuller, 2012; Sontag, 2007), dynamic neural networks (Smith, 1991), and social dynamics (Shi et al., 2019).
@@
Weiss and Margaliot (2021a) introduced a generalization of nonlinear cooperative systems called $k$-cooperative systems. A system is $k$-cooperative if its flow maps the set of vectors with up to $k-1$ sign variations to itself. Just like cooperativity, $k$-cooperativity depends on the sign structure of the system Jacobian, and can be inferred even when the exact values of certain parameters are unknown. The analysis of sign structures guaranteeing $k$-cooperativity relies on the theory of compound matrices (see, e.g., Bar-Shalom et al. (2023)), a fundamental tool also for $k$-contractive (Wu et al., 2022a), $\alpha$-contractive (Wu et al., 2022b), and totally positive differential systems (Margaliot and Sontag, 2019). A sign pattern related to $2$-cooperativity appears in the work by Mallet-Paret and Smith (1990) on monotone cyclic feedback systems (see also Ben Avraham et al. (2020); Feng et al. (2021); Wang et al. (2022) and the references therein). $1$-cooperative systems are cooperative systems, whereas $(n-1)$-cooperative systems (where $n$ is the system dimension) are, up to a coordinate transformation, competitive systems (Weiss and Margaliot, 2021a). Strongly $2$-cooperative systems satisfy a strong Poincaré-Bendixson property; hence, if the closure of a positive orbit corresponding to a bounded solution contains no equilibrium points, then the solution converges to a periodic orbit (Weiss and Margaliot, 2021a).
@@
Bounded solutions that converge not to an equilibrium but to a periodic orbit are of particular interest. For example, in the context of systems biology, Goodwin (1965) states: “It is of fundamental importance to an understanding of cellular organization whether or not the dynamic activity of molecular control processes involves oscillatory behavior.” The theory of cooperative systems cannot support this analysis, because the existence of an attracting periodic orbit automatically implies that the system is not cooperative.
@@
Here, we consider 3-dimensional 2-cooperative systems. Our main result is a new simple sufficient condition for the existence of a set of positive measure such that any bounded solution emanating from this set converges to a periodic orbit, and an explicit characterization of such a set. The analysis is based on combining: (1) the strong Poincaré-Bendixson property of strongly $2$-cooperative systems; (2) the spectral properties of Jacobians of $2$-cooperative systems; and (3) an idea of Rauch (1950) for constructing an invariant set $\Omega$ for a specific 3D model of a non-linear electric circuit so that $\Omega$ does not include an equilibrium. We demonstrate our theoretical results on two well-known non-linear 3D systems in biology: a 3D Goodwin model, and the Field–Noyes ODE model for the famous Belousov–Zhabotinskii reaction.
2 Notation and Preliminary Results
@@
We denote vectors and matrices by lowercase and uppercase letters, respectively. For two vectors $x,y\in{\mathbb{R}}^{n}$, we write $x\geq y$ if $x_{i}\geq y_{i}$ for all $i$. The non-negative orthant in ${\mathbb{R}}^{n}$ is ${\mathbb{R}}^{n}_{\geq 0}:=\{x\in{\mathbb{R}}^{n}\,|\,x\geq 0\}.$ The transpose and determinant of a matrix $A$ are denoted by $A^{\top}$ and $\det(A)$ respectively, while $I_{n}$ denotes the $n\times n$ identity matrix. A square matrix $A$ is Hurwitz if all its eigenvalues have a negative real part, unstable if it has an eigenvalue with a positive real part, and Metzler if all its off-diagonal entries are non-negative. A matrix $\bar{A}$ is a sign pattern matrix (or sign matrix) (Cavers et al., 2020) if every entry of $\bar{A}$ is either * (“don’t care”), 00, $\geq 0$, or $\leq 0$; $\bar{A}$ can also be interpreted as a matrix set. A time-varying matrix $A(t)$ has the sign pattern $\bar{A}$ if the following three properties hold at all times $t$:
@@
    $a_{ij}(t)=0$ for all indices $i,j$ such that $\bar{a}_{ij}$ is 00,
@@
    $a_{ij}(t)\geq 0$ for all indices $i,j$ such that $\bar{a}_{ij}$ is $\geq 0$,
@@
    $a_{ij}(t)\leq 0$ for all indices $i,j$ such that $\bar{a}_{ij}$ is $\leq 0$,
with no restriction on $a_{ij}(t)$ when $\bar{a}_{ij}$ is *. Given a set $S$, $\operatorname{int}(S)$ denotes its interior, and $|S|$ denotes its cardinality.
The flow of a cooperative system maps ${\mathbb{R}}^{n}_{\geq 0}$ to ${\mathbb{R}}^{n}_{\geq 0}$, and it also maps ${\mathbb{R}}^{n}_{\leq 0}:=-({\mathbb{R}}^{n}_{\geq 0})$ to ${\mathbb{R}}^{n}_{\leq 0}$. In other words, the flow maps the set of vectors with zero sign variations to itself. A $k$-cooperative system maps the set of vectors with up to $k-1$ sign variations to itself.
Sign variations in a vector. Let $\sigma(x)$ denote the number of sign variations in a vector $x\in{\mathbb{R}}^{n}$ with no zero entries; for example, $\sigma\left(\begin{bmatrix}6.3&-\pi&1\end{bmatrix}^{\top}\right)=2$. The theory of totally positive matrices (see, e.g., Fallat and Johnson (2011); Gantmacher and Krein (2002); Pinkus (2010)) offers two useful generalizations of $\sigma(\cdot)$ to vectors that may include zero entries.
@@
For any $x\in{\mathbb{R}}^{n}\setminus\{0\}$, let
$s^{-}(x):=\sigma(y),\quad s^{+}(x):=\max_{z\in\mathcal{S}_{x}}\sigma(z),$
where $y$ is the vector obtained from $x$ by deleting all its zero entries, and $\mathcal{S}_{x}$ is the set of vectors obtained from $x$ by replacing each zero entry by either $+1$ or $-1$.
For example, for $x=[-1~{}0~{}0~{}1~{}2]^{\top}$, we have $s^{-}(x)=\sigma([-1~{}1~{}2]^{\top})=1$, whereas $s^{+}(x)=\sigma([-1~{}1~{}-1~{}1~{}2]^{\top})=3$. If $x\in{\mathbb{R}}^{n}$ has no zero entries, then $s^{-}(x)=s^{+}(x)=\sigma(x)$. For the zero vector $0\in{\mathbb{R}}^{n}$, we define $s^{-}(0):=0$, and $s^{+}(0):=n-1$. Then
@@
For example, $P^{1}_{-}={\mathbb{R}}^{n}_{\geq 0}\cup{\mathbb{R}}^{n}_{\leq 0}$, and $P^{1}_{+}=\operatorname{int}({\mathbb{R}}^{n}_{\geq 0}\cup{\mathbb{R}}^{n}_{\leq 0})$. More generally, it can be shown (Weiss and Margaliot, 2021a) that $P^{k}_{-}$ is closed, and that $P^{k}_{+}=\operatorname{int}(P^{k}_{-})$ for all $k$.
A set $C\subseteq{\mathbb{R}}^{n}$ is a cone if $x\in C$ implies that $\alpha x\in C$ for all $\alpha\in{\mathbb{R}}_{\geq 0}$. The set $P^{k}_{-}$ is a cone for all $k$. However, it is not a convex cone. For example, given vectors $y_{1}:=[-2~{}1~{}1]^{\top}\in P^{2}_{-}$ and $y_{2}:=[1~{}1~{}-2]^{\top}\in P^{2}_{-}$, their convex combination $\frac{1}{2}y_{1}+\frac{1}{2}y_{2}=\frac{1}{2}[-1~{}2~{}-1]^{\top}\not\in P^{2}_{-}$.
To understand the geometry of $P^{k}_{-}$, recall that a closed set $C\subseteq{\mathbb{R}}^{n}$ is a cone of rank $k$ if (1) $x\in C$ implies that $\alpha x\in C$ for all $\alpha\in{\mathbb{R}}$ (note that the requirement is for all $\alpha\in{\mathbb{R}}$, not just for all $\alpha\in{\mathbb{R}}_{\geq 0}$); and (2) $C$ contains a linear space of dimension $k$, and no linear space of a higher dimension (Krasnoselśkii et al., 1989). The set $P^{1}_{-}={\mathbb{R}}^{n}_{\geq 0}\cup{\mathbb{R}}^{n}_{\leq 0}$ is a cone of rank $1$. A cone $C$ of rank $k$ is $k$-solid if there exists a $k$-dimensional linear subspace $V$ such that $V\setminus\{0\}\subseteq\operatorname{int}(C)$. For all $k\in\{1,\dots,n\}$, the set $P^{k}_{-}$ is a cone of rank $k$ that is $k$-solid (Weiss and Margaliot, 2021a).
Roughly speaking, if a dynamical system admits a $k$-solid cone $C$ as an invariant set, and if its trajectories can be projected in a one-to-one way on the $k$-dimensional linear subspace $V$, then its trajectories “behave” as those of a $k$-dimensional system. In particular, if $k=2$, then the trajectories “behave” as those of a planar dynamical system (Sanchez, 2009).
@@
Assume that $t\to A(t)$ is continuous. Then, the linear system (1) is $1$-positive (i.e., positive) iff $A(t)$ is Metzler for all $t$. The linear system (1) is $2$-positive iff, for all $t$$A(t)$ has the sign pattern
$\bar{A}_{2}:=\begin{bmatrix}*&\geq 0&0&\dots&0&\leq 0\\ \geq 0&*&\geq 0&\dots&0&0\\ 0&\geq 0&*&\dots&0&0\\ \vdots&\vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&0&\dots&*&\geq 0\\ \leq 0&0&0&\dots&\geq 0&*\end{bmatrix},$ (2)
while it is strongly $2$-positive if, in addition, the matrix $A(t)$ is irreducible for almost all $t$ (Weiss and Margaliot, 2021a) (see also Ge and Arcak (2009)). Since the upper-right and lower-left corner entries of $A(t)$ may take negative values, $A(t)$ is not necessarily Metzler.
$k$-cooperative nonlinear dynamical systems. A dynamical system is $k$-cooperative if its associated variational system, which is an LTV system (see, e.g., Wu et al. (2022a)), is $k$-positive for all $t\geq 0$ and all $x$ in the state space. Consider the non-linear time-invariant system
$\dot{x}(t)=f(x(t)),$ (3)
@@
$\dot{z}(t)=M(t)z(t),$ (4)
where $M(t):=\int_{0}^{1}J\big{(}rx(t,a)+(1-r)x(t,b)\big{)}\mathop{}\!\mathrm{d}r$.
System (4) is LTV and, since sign patterns are preserved under integration, if $J(x)$ has some sign pattern for all $x\in\Omega$, then $M(t)$ has the same sign pattern for all $t\geq 0$.
@@
For example, the nonlinear system (3) is $1$-cooperative (i.e., cooperative) if the variational system (4) is $1$-positive, that is, if $J(x)$ is Metzler for all $x$.
For an initial condition $a\in\Omega$, let $\omega(a)$ denote the omega limit set of the solution of (3) emanating from $x(0)=a$, namely, the set of all points $y\in\Omega$ for which there exists a sequence $(t_{n})_{n\in\mathbb{N}}$ in ${\mathbb{R}}$, with $\lim_{n\to\infty}t_{n}=+\infty$, such that $\lim_{n\to\infty}x(t_{n},a)=y$; see e.g. (Teschl, 2012, p. 193).